# The Narrow Gate: Localized Image-Text Communication in Native Multimodal Models

### v3

@@ 1 Introduction
Although recent studies have examined the internal mechanisms of unimodal output VLMs [27, 28, 29, 30], it remains largely unknown how modalities interact in the hidden representations of native multimodal models. In this work, we focus on some of the latest native VLMs, Chameleon [20] and Emu3 [26], and compare them with a range of non-native ones, including both text-only VLMs such as LLaVA [11] and multimodal generators that rely on a pretrained language backbone like VILA-U [31], on how they transfer information from the visual domain to the textual domain in various image understanding tasks.
We observe that in native multimodal models (Chameleon and Emu3), image and text representations remain well-separated in different clusters from input to output, while for non-native models they tend to mix in late layers (section˜3.1). Additionally, in this second class of models, visual information flows to text through multiple tokens. In contrast, native multimodal VLMs channel the global image information into a single token placed immediately after the image, the end-of-image token, denoted by [EOI](section˜3.2). The [EOI]token behaves like a memory token [32] or narrow gate through which the image information must pass to guide text generation.
@@ 2.1 Analytical Tools
We used the HuggingFace implementations Chameleon [46, 47], Emu3-Gen [48], LLaVA [49], Pixtral [50], VILA-U [31] and Janus [51] models. We run all the experiments on a single NVIDIA A100 GPUs with 40GB VRAM. For the Chameleon 34B, we used two 40GB GPUs. The code needed to reproduce the experiments is available at: ritareasciencepark.github.io/Narrow-gate.
@@ 3.1 Modality Gap in Native Multimodal VLMs
To address this question, we randomly select $10,000$ image-caption pairs from the Flickr30k dataset and extract the residual stream representations of all hidden layers at a randomly chosen text token position (see appendix˜B for the complete prompt structure). To compare the geometric organization of the different modalities, we first measure the median cosine similarity between representations of image and text at each hidden layer.
As shown in LABEL:fig:modality-gap-left, this analysis reveals different trends across architectures. In the Chameleon-7B (blue) and Chameleon-34B (green), the representative vectors of image and text tokens remain nearly orthogonal throughout the hidden layers, with median cosine similarity values consistently below $0.10$.
@@ 3.3 Probing the Semantic Content of Visual Tokens
A similar trend is observed in other non-native multimodal, as is shown in figure˜A3 models such as Janus, Pixtral, and VILA-U: internal image tokens retain the highest $\chi_{k}^{l,gt}$, while [EOI]contributes little semantic signal. These results suggest two distinct encoding regimes. In native multimodal models (Chameleon, Emu3), the [EOI]token both attracts the majority of text-on-image attention and carries the richest visual semantics, fulfilling the conditions for a narrow communication gate. In non-native multimodal models, semantic content remains broadly distributed across internal image tokens, indicating a diffuse, multi-token communication pattern.
@@ 3.4 Ablation Experiments: The Effect of Localized Communication on Downstream Tasks
For each model, we randomly sample 2,0002{,}000 examples per dataset and measure baseline performance before applying targeted attention knockouts. The performance for visual question answering VQAv2 are evaluated with the VQA metric [54] , for the captioning benchamarks instead we used CIDEr [55]. We block (i) communication between the [EOI] and text tokens by zeroing out $A_{\text{text}\rightarrow\texttt{[EOI]}}$, and (ii) all text-on-image attention by zeroing out $A_{\text{text}\rightarrow\text{img}}$ (section˜2.1). This isolates the impact of the [EOI] token relative to the broader visual stream. Results for Chameleon, Emu3, and LLaVA are summarized in table˜1; additional models (Pixtral, Janus, VILA-U) and further token positions are reported in section˜C.4.
@@ 4.2 Improving Model Robustness with Localized Fine-tuning
The presence of a narrow gate in native multimodal models can constrain how the model integrate visual and linguistic features, making its representations overly dependent on a specific internal pathway. To promote a richer and more distributed exchange between modalities, in this section, we show a fine-tuning strategy that masks the [EOI]token during training, forcing the model to encode and communicate visual information stored in [EOI]through alternative internal tokens. Details about the training setup are reported in section˜B.3.
@@ 5 Discussion and Conclusion
The presence of a narrow gate is linked to the modality gap between image and text representations. Based on our results we hypothesized that three architectural and training factors jointly influence its emergence. The multimodal output objective: a training objective that requires generating both image and text tokens, encourages the model to maintain distinct representational pathways for each modality, enforcing their separation. The training history: models trained from scratch on multimodal data, develop different subspaces for each modality subspaces, while models derived from pre-trained text-only LLMs and later fine-tuned for multimodal use (Janus, VILA-U) retain a text-aligned internal geometry and exhibit distributed communication. The type of image encoder: low-level visual tokenizers such as VQ-GAN (used in Chameleon), represent images through local features rather than high-level semantic embeddings. This increases the abstraction gap between image and text tokens and promotes the formation of a distinct communication interface centered on [EOI]. Taken together, these factors show that the narrow gate is not an incidental artifact but a structural consequence of jointly training a multimodal generator with low-level image tokenization.
@@ Appendix A Related Works
Information Flow in Text-Only VLMs.
In text-only VLMs, this kind of summarization is particularly useful since much of the attention to image tokens is concentrated in the initial layers, likely on a few “anchor tokens” [28, 59]. Basu et al. [28] demonstrated with multimodal causal tracing that, in VQA tasks, a subset of strongly attended late image tokens in early layers transfer information to the text. However, a later ablation study by Neo et al. [27] showed that important information about objects is often localized in tokens corresponding to the object’s position in the image. Interestingly, they also observed that image tokens, which may have different statistical properties compared to textual tokens [60], in the deeper layers encode vocabulary words that describe the objects they belong to.
@@ B.2 Datasets details
The dataset VQAv2 [37] consists of $204,721$ images from the MSCOCO dataset each associated with a variable number multiple choice question answering. For this benchmark presented in tables˜1 and A3 we constructed prompts concatenating the image, the first of the associated question, and the instruction “Answer the question using a single word or phrase.
@@
For the activation patching experiment, described in section˜4.1, we manually select $20$ class from the list above to ensure semantic diversity in the animal represented, and then we random sample $100$ images for each class, obtaining a total of $2000$ images. The selected pairs are:
@@ C.1 Modality Gap in VLMs
We extend the results of section˜3.1 to Janus, Pixtral and VILA-U. As illustrated in Figure˜A1, all of the additional models exhibit an alignment between modalities that grows along the model depth, similar to the behavior observed in LLaVA. Figure˜A1-left confirms this alignment across modalities for all the models. Figure˜A1-right reveals that the homogeneity within modality clusters is lower compared to the Chameleon models and Emu3. This indicates that LLaVA, Janus, Pixtral and VILA-U have a smaller modality gap.
Figure A1: Modality Gap in VLMs. (left) Cosine similarity between text and image token embeddings as a function of model depth reflects the orthogonality of modalities in Janus, Pixtral and VILA-U models. Points represent median cosine similarity, with shaded areas indicating the interquartile range. (right) Homogeneity score of token clusters generated via Advanced Density Peaks with respect to their original modality.
@@
In section˜3.3 we investigated the alignment between token representations across layers and the classes of ImageNet. In this section, we will now extend those findings to the MS-COCO captioning task to investigate if the results generalize to tasks different from classification. We will proceed as follows: first, for each image in the MS-COCO dataset, we extract the activations of the relevant visual tokens at each layer of every model analyzed in our study; Concurrently, we embedded the captions for these images using the Qwen2-7B-GTE text encoder [73]. To measure the alignment between these two sets of representations we used neighborhood overlap described at section˜2.1 where in this context the reference data representations are the embeddings from the text encoder. In other words for each sample $i$, we computed the fraction of shared $k$-nearest neighbors between the visual token embedding and the caption embedding according to equation˜2. This approach allowed us to perform a direct comparison of the semantic alignment across all models and token types.
@@ C.3 Cross-Modal Attention and Semantic Content of Visual Tokens
The results for Chameleon-34B closely mirror those of Chameleon-7B and Emu3, as discussed in section˜3.2. Figure˜A3-left visualizes the cross-modal attention contributions of image tokens, where tokens contributing more than $1\%$ individually are shown explicitly, while the rest are grouped as "internal image". In Chameleon-34B, cross-modal attention is primarily concentrated on three tokens—[EOI], the first image token, and the last. In contrast, Janus, Pixtral, and VILA-U exhibit a pattern similar to LLaVA (section˜3.2), where attention to special tokens remains low. At the same time, a significant portion of the information is retained within internal image tokens. Figure˜A3-center further reveals that in Chameleon-34B as in the other native multimodal VLMs, [EOI]consistently maintains high $\chi_{k}^{l,gt}$ values across layers, indicating its dominant role in encoding visual information. Meanwhile, all other tokens exhibit near-zero overlap values, suggesting minimal contribution to meaningful visual representation. For Janus, both [EOI]and internal image tokens carry strong information from the earliest layers, implying that visual encoding may be directly influencing their representations. Similarly in Pixtral, both the [EOI]and the last image token exhibit a $\chi_{k}^{l,gt}$ higher than that of LLaVA, although still below Janus. However, as shown in section˜C.4, the information encoded in these special tokens do not play a significant role in information flow within the backbone models. Among the non-native multimodal VLMs, VILA-U most closely resembles LLaVA in terms of semantic alignment, as most of the information is concentrated within the internal image tokens, while all other special tokens achieving negligible overlap scores.
@@ C.4 Effect of attention knockout on vision-language tasks
Table˜A3 extends the results of table˜1, showing the ablation effect on all tokens identified for their high cross-modal attention. The analysis further confirms that none of the special tokens, except for [EOI], play a significant role, as their ablation does not noticeably affect performance on any benchmarks.

### v4

@@ 1 Introduction
Although recent studies have examined the internal mechanisms of unimodal output VLMs [45, 12, 67, 62], it remains largely unknown how modalities interact in the hidden representations of native multimodal models. In this work, we focus on some of the latest native VLMs, Chameleon [57] and Emu3 [66], and compare them with a range of non-native ones, including both text-only VLMs such as LLaVA [44] and multimodal generators that rely on a pretrained language backbone like VILA-U [72], on how they transfer information from the visual domain to the textual domain in various image understanding tasks. We observe that in native multimodal models (Chameleon and Emu3), image and text representations remain well-separated in different clusters from input to output, while for non-native models they tend to mix in late layers (section˜3.1). Additionally, in this second class of models, visual information flows to text through multiple tokens. In contrast, native multimodal VLMs channel the global image information into a single token placed immediately after the image, the end-of-image token, denoted by [EOI](section˜3.2). The [EOI]token behaves like a memory token [16] or narrow gate through which the image information must pass to guide text generation.
@@ 2.1 Analytical Tools
We used the HuggingFace implementations of Chameleon [4, 3], Emu3-Gen [9], LLaVA [59], Pixtral [5], VILA-U [72] and Janus [2] models. We run all the experiments on a single NVIDIA A100 GPUs with 40GB VRAM. For the Chameleon 34B, we used two 40GB GPUs. The code needed to reproduce the experiments is available at: ritareasciencepark.github.io/Narrow-gate.
@@ 3.1 Modality Gap in Native Multimodal VLMs
To address this question, we randomly select $10,000$ image-caption pairs from the Flickr30k dataset and extract the residual stream representations of all hidden layers at a randomly chosen text token position (see appendix˜B for the complete prompt structure). To compare the geometric organization of the different modalities, we first measure the median cosine similarity between representations of image and text at each hidden layer. As shown in LABEL:fig:modality-gap-left, this analysis reveals different trends across architectures. In the Chameleon-7B (blue) and Chameleon-34B (green), the representative vectors of image and text tokens remain nearly orthogonal throughout the hidden layers, with median cosine similarity values consistently below $0.10$.
@@ 3.3 Probing the Semantic Content of Visual Tokens
A similar trend is observed in other non-native multimodal models, as is shown in figure˜A3 models such as Janus, Pixtral, and VILA-U: internal image tokens retain the highest $\chi_{k}^{l,gt}$, while [EOI]contributes little semantic signal. These results suggest two distinct encoding regimes. In native multimodal models (Chameleon, Emu3), the [EOI]token both attracts the majority of text-on-image attention and carries the richest visual semantics, fulfilling the conditions for a narrow communication gate. In non-native multimodal models, semantic content remains broadly distributed across internal image tokens, indicating a diffuse, multi-token communication pattern.
@@ 3.4 Ablation Experiments: The Effect of Localized Communication on Downstream Tasks
For each model, we randomly sample 2,0002{,}000 examples per dataset and measure baseline performance before applying targeted attention knockouts. The performance for visual question answering VQAv2 is evaluated with the VQA metric [8], for the captioning benchamarks instead we used CIDEr [65]. We block (i) communication between the [EOI] and text tokens by zeroing out $A_{\text{text}\rightarrow\texttt{[EOI]}}$, and (ii) all text-on-image attention by zeroing out $A_{\text{text}\rightarrow\text{img}}$ (section˜2.1). This isolates the impact of the [EOI] token relative to the broader visual stream. Results for Chameleon, Emu3, and LLaVA are summarized in table˜1; additional models (Pixtral, Janus, VILA-U) and further token positions are reported in section˜C.4.
@@ 4.2 Improving Model Robustness with Localized Fine-tuning
The presence of a narrow gate in native multimodal models can constrain how the model integrates visual and linguistic features, making its representations overly dependent on a specific internal pathway. To promote a richer and more distributed exchange between modalities, in this section, we show a fine-tuning strategy that masks the [EOI]token during training, forcing the model to encode and communicate visual information stored in [EOI]through alternative internal tokens. Details about the training setup are reported in section˜B.3.
@@ 5 Discussion and Conclusion
The presence of a narrow gate is linked to the modality gap between image and text representations. Based on our results, we hypothesized that three architectural and training factors jointly influence its emergence. The multimodal output objective: a training objective that requires generating both image and text tokens, encourages the model to maintain distinct representational pathways for each modality, enforcing their separation. The training history: models trained from scratch on multimodal data, develop different subspaces for each modality, while models derived from pre-trained text-only LLMs and later fine-tuned for multimodal use (Janus, VILA-U) retain a text-aligned internal geometry and exhibit distributed communication. The type of image encoder: low-level visual tokenizers such as VQ-GAN (used in Chameleon), represent images through local features rather than high-level semantic embeddings. This increases the abstraction gap between image and text tokens and promotes the formation of a distinct communication interface centered on [EOI]. Taken together, these factors show that the narrow gate is not an incidental artifact but a structural consequence of jointly training a multimodal generator with low-level image tokenization.
@@ Appendix A Related Works
Localization of information in Text-Only VLMs.
In text-only VLMs, this kind of summarization is particularly useful since much of the attention to image tokens is concentrated in the initial layers, likely on a few “anchor tokens” [12, 19]. Basu et al. [12] demonstrated with multimodal causal tracing that, in VQA tasks, a subset of strongly attended late image tokens in early layers transfer information to the text. However, subsequent studies have shown that important visual information is localized in heads of deeper layers of the network [11, 46], or within tokens corresponding to the spatial positions of objects in the image [45]. Interestingly, although these tokens have different statistical properties from those of textual tokens [17], they encode, in deeper layers, vocabulary words that describe the objects to which they correspond.
@@ B.2 Datasets details
The dataset VQAv2 [34] consists of $204,721$ images from the MSCOCO dataset each associated with a variable number multiple choice question answering. For this benchmark presented in tables˜1 and A3, we constructed prompts concatenating the image, the first of the associated question, and the instruction “Answer the question using a single word or phrase.
@@
For the activation patching experiment, described in section˜4.1, we manually select $20$ classes from the list above to ensure semantic diversity in the animal represented, and then we random sample $100$ images for each class, obtaining a total of $2000$ images. The selected pairs are:
@@ C.1 Modality Gap in VLMs
We extend the results of section˜3.1 to Janus, Pixtral, and VILA-U. As illustrated in Figure˜A1, all of the additional models exhibit an alignment between modalities that grows along the model depth, similar to the behavior observed in LLaVA. Figure˜A1-left confirms this alignment across modalities for all the models. Figure˜A1-right reveals that the homogeneity within modality clusters is lower compared to the Chameleon models and Emu3. This indicates that LLaVA, Janus, Pixtral, and VILA-U have a smaller modality gap.
Figure A1: Modality Gap in VLMs. (left) Cosine similarity between text and image token embeddings as a function of model depth reflects the orthogonality of modalities in Janus, Pixtral, and VILA-U models. Points represent median cosine similarity, with shaded areas indicating the interquartile range. (right) Homogeneity score of token clusters generated via Advanced Density Peaks with respect to their original modality.
@@
In section˜3.3 we investigated the alignment between token representations across layers and the classes of ImageNet. In this section, we will now extend those findings to the MS-COCO captioning task to investigate if the results generalize to tasks different from classification. We will proceed as follows: first, for each image in the MS-COCO dataset, we extract the activations of the relevant visual tokens at each layer of every model analyzed in our study; Concurrently, we embedded the captions for these images using the Qwen2-7B-GTE text encoder [40]. To measure the alignment between these two sets of representations we used neighborhood overlap described at section˜2.1, where in this context, the reference data representations are the embeddings from the text encoder. In other words for each sample $i$, we computed the fraction of shared $k$-nearest neighbors between the visual token embedding and the caption embedding according to equation˜2. This approach allowed us to perform a direct comparison of the semantic alignment across all models and token types.
@@ C.3 Cross-Modal Attention and Semantic Content of Visual Tokens
The results for Chameleon-34B closely mirror those of Chameleon-7B and Emu3, as discussed in section˜3.2. Figure˜A3-left visualizes the cross-modal attention contributions of image tokens, where tokens contributing more than $1\%$ individually are shown explicitly, while the rest are grouped as "internal image". In Chameleon-34B, cross-modal attention is primarily concentrated on three tokens—[EOI], the first image token, and the last. In contrast, Janus, Pixtral, and VILA-U exhibit a pattern similar to LLaVA (section˜3.2), where attention to special tokens remains low. At the same time, a significant portion of the information is retained within internal image tokens. Figure˜A3-center further reveals that in Chameleon-34B as in the other native multimodal VLMs, [EOI]consistently maintains high $\chi_{k}^{l,gt}$ values across layers, indicating its dominant role in encoding visual information. Meanwhile, all other tokens exhibit near-zero overlap values, suggesting minimal contribution to meaningful visual representation. For Janus, both [EOI]and internal image tokens carry strong information from the earliest layers, implying that visual encoding may be directly influencing their representations. Similarly in Pixtral, both the [EOI]and the last image token exhibit a $\chi_{k}^{l,gt}$ higher than that of LLaVA, although still below Janus. However, as shown in section˜C.4, the information encoded in these special tokens do not play a significant role in information flow within the backbone models. Among the non-native multimodal VLMs, VILA-U most closely resembles LLaVA in terms of semantic alignment, as most of the information is concentrated within the internal image tokens, while all other special tokens achieve negligible overlap scores.
@@ C.4 Effect of attention knockout on vision-language tasks
Table˜A3 extends the results of table˜1, showing the ablation effect on all tokens identified for their strong cross-modal attention. The analysis further confirms that none of the special tokens, except for [EOI], play a significant role, as their ablation does not noticeably affect performance on any benchmarks.